# LIT-LVM: Structured Regularization for Interaction Terms in Linear Predictors using Latent Variable Models

### v1

@@ Abstract
Some of the simplest, yet most frequently used predictors in statistics and machine learning use weighted linear combinations of features. Such linear predictors can model non-linear relationships between features by adding interaction terms corresponding to the products of all pairs of features. We consider the problem of accurately estimating coefficients for interaction terms in linear predictors. We hypothesize that the coefficients for different interaction terms have an approximate low-dimensional structure and represent each feature by a latent vector in a low-dimensional space. This low-dimensional representation can be viewed as a structured regularization approach that further mitigates overfitting in high-dimensional settings beyond standard regularizers such as the lasso and elastic net. We demonstrate that our approach, called LIT-LVM, achieves superior prediction accuracy compared to elastic net and factorization machines on a wide variety of simulated and real data, particularly when the number of interaction terms is high compared to the number of samples. LIT-LVM also provides low-dimensional latent representations for features that are useful for visualizing and analyzing their relationships.
@@ 1 Introduction
Linear predictors such as linear regression and logistic regression model an example’s target variable $y$ using a weighted linear combination of its features $x_{1},x_{2},\ldots,x_{p}$. These models are favored for their speed, simplicity, and interpretability. In high-dimensional settings where the number of features $p$ is close to or even exceeds the number of examples $n$, regularization techniques such as the lasso Tibshirani (1996) and elastic net Zou & Hastie (2005), have become essential to prevent overfitting and enhance prediction accuracy on unseen test data.
@@
Unlike the original features, interaction terms have additional structure, as they denote relationships between two features. The coefficients for these interaction terms can be arranged in a $p\times p$ coefficient matrix $\mathbf{\bm{\Theta}}$, where entry $\theta_{jk}$ denotes the coefficient for the interaction term $x_{j}x_{k}$. We hypothesize that this matrix has an approximate low-dimensional structure, so that we can represent each feature $j$ by a vector $\mathbf{\bm{z}}_{j}$ in a $d$-dimensional space, where $d<p$.
@@
Our main contribution is a structured regularization approach for estimating the coefficients for interaction terms in linear predictors using an approximate low-dimensional structure, as shown in Figure 1, in addition to traditional regularization techniques such as the elastic net. We call our proposed approach LIT-LVM, denoting linear interaction terms with a latent variable model, and is applicable to many types of linear predictors, including linear regression, logistic regression, and the Cox proportional hazards (Cox PH) model (Cox, 1972).
@@ 2.1 Linear Predictors
$\hat{y}=f(\mathbf{\bm{\beta}}^{T}\mathbf{\bm{x}}+\mathbf{\bm{\theta}}_{\text{flat}}^{T}\mathbf{\bm{x}}_{\text{int}})=f(\tilde{\mathbf{\bm{\beta}}}^{T}\tilde{\mathbf{\bm{x}}}),$ (3)
@@ 2.2.2 Latent Variable Models
2.2.3 Factorization Machines (FMs)
The first FM model was proposed by Rendle (2010). It assumes that each feature $j$ has a $d$-dimensional latent vector $\mathbf{\bm{z}}_{j}$ such that so that $\mathbf{\bm{\Theta}}=\mathbf{\bm{Z}}\mathbf{\bm{Z}}^{T}$. This forces $\mathbf{\bm{\Theta}}$ to have an exact low-rank structure. FMs have also been extended to higher-order interaction terms in a scalable manner Blondel et al. (2016).
FMs differ from our proposed LIT-LVM approach in two main aspects. First, they are often applied to data with both large $n$ and $p$ with a sparse feature matrix $\mathbf{\bm{X}}$ and are heavily concerned with scalability when using interaction terms Rendle (2012; 2013). On the other hand, our focus is on improving accuracy of estimation and prediction, not scalability. Second, FMs assume that $\mathbf{\bm{\Theta}}$ is exactly low rank, whereas we assume that $\mathbf{\bm{\Theta}}$ has an approximate low-dimensional structure (possibly low rank, but not necessarily). This is a subtle, yet important distinction, and we provide more details in Section 3.3.
@@ 3.1 LVM for Interaction Terms
where the intercept $\alpha_{0}$ denotes the baseline interaction level.
@@ 3.2 Estimation Procedure
We use Adam Kingma (2014) with various learning rates $\in$ [0.005,0.01,0.05,0.1[0.005,0.01,0.05,0.1[ 0.005 , 0.01 , 0.05 , 0.1] as our optimizer for the total loss function. To handle the $\ell_{1}$ penalty, we use the proximal gradient method Parikh et al. (2014), where the proximal operator corresponds to the soft thresholding operator. The other terms are all differentiable, with derivatives computed by automatic differentiation in PyTorch. The initializations of $\tilde{\mathbf{\bm{\beta}}}$ and $\mathbf{\bm{Z}}$ are drawn from normal distributions: $\tilde{\beta}_{j}\sim\mathcal{N}(0,1)$ and $z_{jk}\sim\mathcal{N}(0,1)$. The time complexity of our estimation procedure per epoch is $O(np^{2})$, as we discuss in Section A.2.
@@ 4 Simulation Experiments
We begin with some simulation experiments to investigate the effects of LIT-LVM as we change the number of features $p$ with fixed number of examples $n$. As $p$ increases, the model becomes more difficult to estimate. Code to reproduce all experiments is available at https://github.com/MLNS-Lab/LIT-LVM/.
@@ Results
Figure 2: Results of simulation experiments on linear and logistic regression where the true latent dimension for the interaction weights $\mathbf{\bm{\Theta}}$ is $d_{true}=2$. (2(a)) Lower RMSE and (2(b)) higher AUC denote higher prediction accuracy. Error bars denote one standard error. Our proposed LIT-LVM approach, which uses structured regularization, significantly outperforms the elastic net with interaction terms, a traditional regularization approach, as $p$ increases for fixed $n$.
The results for fixed $n=1,\!000$, true latent dimension $d_{\text{true}}=2$, and four different values of $p$ are shown in Figure 2(a). We compare our proposed LIT-LVM approach using a low rank model to elastic net with interaction terms. For $p=40$, $\mathbf{\bm{\Theta}}$ has $\binom{40}{2}=780$ entries, which is less than $n$, and both approaches are quite accurate. At $p=50$, $\mathbf{\bm{\Theta}}$ has $\binom{50}{2}=1,225$ entries, which now exceeds $n$, and we observe a phase transition where the RMSE of elastic net increases significantly compared to our LIT-LVM approach. As $p$ continues to increase, LIT-LVM continues to outperform the elastic net with interactions.
@@ 4.2 Logistic Regression with Latent Distance
In this experiment, we generate the feature matrix and interaction terms in the same manner as in the linear regression experiment in Section 4.1. For logistic regression, each element of the response vector $\mathbf{\bm{y}}$ is sampled from a Bernoulli distribution with success probability given by $\operatorname{logistic}(\mathbf{\bm{\beta}}^{T}\mathbf{\bm{x}}+\mathbf{\bm{\theta}}_{\text{flat}}^{T}\mathbf{\bm{x}}_{\text{int}})+\eta_{i})$, where $\operatorname{logistic}(x)=\frac{1}{1+e^{-x}}$. We inject sparsity into the augmented coefficient vector $\tilde{\mathbf{\bm{\beta}}}$ by first generating a dense version $\tilde{\mathbf{\bm{\beta}}}^{\text{dense}}$. We then set each element $\tilde{\beta}^{\text{dense}}_{j}$ to 0 with probability given by the Gaussian function $\exp(-(\tilde{\beta}^{\text{dense}}_{j})^{2}/\sigma_{s}^{2})$, where $\sigma_{s}^{2}$ is a parameter that adjusts the level of sparsity. This ensures that larger elements in $\tilde{\mathbf{\bm{\beta}}}^{\text{dense}}$ have a lower probability of being zeroed out.
@@ Results
The results for fixed $n=1,\!000$, true latent dimension $d_{\text{true}}=2$, and five different values of $p$ are shown in Figure 2(b). Similar to the results from the linear regression experiment, we observe a phase transition where the accuracy of elastic net decreases significantly compared to our approach. The phase transition occurs at roughly $p=30$, lower than in the linear regression experiment. As $p$ continues to increase, our LIT-LVM approach continues to outperform elastic net, with prediction accuracy decreasing much slower than in elastic net.
@@
Comparison with FMs
In Section B.3, we conduct simulation experiments comparing the performance of LIT-LVM with FMs. We simulate data in the same manner except with a low rank structure rather than a latent space structure, to better match the assumption of FMs. Likewise, we use LIT-LVM with a low rank model. We find that FMs perform worse when $\mathbf{\bm{\Theta}}$ is noisy or exhibits higher sparsity (deviation from exact low rank). This indicates that factorizing the interaction coefficients, as FMs do, can degrade performance when the underlying assumption of exact low rank structure is not met. In contrast, the approximate low rank structure assumption in LIT-LVM offers greater flexibility in modeling the interaction coefficients, ensuring more robustness and higher accuracy.
@@
Figure 3: Mean AUC for different values of the LVM regularization strength $\lambda_{l}$ on simulated data experiments using logistic regression. (3(a)) The simulation with low noise ($\theta^{2}_{\theta}=0.1$) peaks in AUC at an intermediate $\lambda_{l}$, confirming the benefit of the LVM regularization term. (3(b)) The simulation with high noise ($\theta^{2}_{\theta}=4$) has destroyed the latent structure, and thus shows minimal benefit from the LVM regularization.
@@
To assess the robustness of LIT-LVM under severe violations of the low-rank assumption, we repeat the same simulation but increase latent variable noise variance to $\sigma_{\theta}^{2}=4$, effectively destroying the underlying low-rank structure. As shown in Figure 3(b), the best performing $\lambda_{l}$ is near 00 or $0.01$ which corresponds to applying no or little latent penalty. This is not surprising because at such high noise, the interaction matrix no longer possesses a low-rank structure so a larger values of $\lambda_{l}$ degrades the performance.
@@ 5.1 OpenML Datasets
We conduct experiments on a wide variety of publicly available datasets from OpenML Vanschoren et al. (2013) to verify how well our structured regularization for the interaction matrix applies to real data. We consider both regression and binary classification tasks using linear and logistic regression, respectively. We consider 12 regression datasets and 10 classification datasets, described in Section C.1. For each task and dataset, we compare the accuracy of four different predictors: the base linear predictor (linear or logistic regression) applied directly to the features, the same linear predictor with interaction terms, our LIT-LVM approach, and factorization machines (FMs). We apply an elastic net penalty for all four methods.
We divide the data into a 50/50 train/test split for both classification and regression tasks. We tune hyperparameters using a grid search, selecting the set of hyperparameters that minimize the mean RMSE for regression tasks and maximize the mean AUC for classification tasks over five splits within the training set. Subsequently, we evaluate prediction accuracy on the held-out test set. We repeat this process using five different train/test splits to calculate the standard errors for the prediction accuracy.
@@ Results
Dataset Name $p^{2}/n$ Elastic Net (No Interactions) Elastic Net (with Interactions) LIT-LVM FM
boston $0.153$ $4.99\pm 0.12$ $\mathbf{4.39\pm 0.23}$ $\mathit{4.40\pm 0.26}$ $5.012\pm 0.63$
bike_sharing_demand $0.0207$ $148.40\pm 1.08$ $\mathit{144.48\pm 1.06}$ $\mathbf{144.47\pm 1.05}$ $156.65\pm 1.073$
diamonds $0.00185$ $0.28\pm 0.002$ $0.26\pm 0.002$ $\mathbf{0.25\pm 0.002}$ $0.283\pm 0.003$
fri_c4_500_10 $0.2$ $0.88\pm 0.014$ $0.89\pm 0.015$ $\mathbf{0.87\pm 0.017}$ $0.997\pm 0.031$
medical_charges $0.000883$ $0.24\pm 0.000$ $\mathbf{0.19\pm 0.003}$ $\mathbf{0.19\pm 0.003}$ $0.204\pm 0.003$
nyc_taxi_green $0.000138$ $0.50\pm 0.001$ $0.50\pm 0.001$ $\mathbf{0.16\pm 0.003}$ $0.50\pm 0.001$
pol $0.0451$ $30.55\pm 0.08$ $\mathit{24.33\pm 0.71}$ $\mathbf{24.24\pm 0.63}$ $29.05\pm 0.331$
rmftsa_ladata $0.196$ $1.86\pm 0.05$ $\mathit{1.73\pm 0.05}$ $\mathbf{1.72\pm 0.03}$ $1.85\pm 0.053$
socmob $1.31$ $23.57\pm 0.61$ $\mathit{17.17\pm 1.15}$ $\mathbf{16.55\pm 1.56}$ $23.56\pm 1.956$
space_ga $0.158$ $0.14\pm 0.004$ $0.13\pm 0.004$ $\mathbf{0.12\pm 0.000}$ $0.134\pm 0.0.004$
superconduct $0.294$ $17.72\pm 0.07$ $15.17\pm 0.19$ $\mathbf{14.32\pm 0.07}$ $17.704\pm 0.082$
wind $0.034$ $3.30\pm 0.014$ $3.28\pm 0.033$ $\mathbf{3.17\pm 0.016}$ $3.277\pm 0.012$
@@
Table 2: Mean AUC $\pm$ standard error over 5 different train/test splits for OpenML classification datasets. Higher AUC denotes higher prediction accuracy. Highest accuracy for each data set is denoted in bold, and entries within 1 standard error are denoted in italic. Our proposed LIT-LVM approach has the highest accuracy on 9 of the 10 datasets and is within 1 standard error on the other 4 datasets.
@@
Dataset Name $p^{2}/n$ Elastic Net (No Interactions) Elastic Net (with Interactions) LIT-LVM FM
Bioresponse $51.4$ $0.791\pm 0.013$ $0.791\pm 0.004$ $\bm{0.808\pm 0.003}$ $0.801\pm 0.004$
clean1 $60.0$ $0.900\pm 0.003$ $0.922\pm 0.004$ $\bm{0.947\pm 0.002}$ $0.900\pm 0.005$
clean2 $4.33$ $0.944\pm 0.001$ $0.943\pm 0.003$ $\bm{0.985\pm 0.001}$ $0.955\pm 0.001$
eye_movements $0.0748$ $0.586\pm 0.004$ $\bm{0.608\pm 0.002}$ $\mathit{0.607\pm 0.002}$ $0.586\pm 0.004$
fri_c4_500_100 $20.4$ $\mathit{0.666\pm 0.011}$ $0.576\pm 0.020$ $\bm{0.666\pm 0.010}$ $0.651\pm 0.012$
fri_c4_1000_100 $10.2$ $\mathit{0.672\pm 0.009}$ $0.564\pm 0.009$ $\bm{0.689\pm 0.008}$ $0.674\pm 0.014$
hill-valley $8.41$ $\mathit{0.711\pm 0.020}$ $0.547\pm 0.027$ $\bm{0.734\pm 0.034}$ $\mathit{0.699\pm 0.023}$
jannis $0.0525$ $0.814\pm 0.000$ $\mathit{0.826\pm 0.001}$ $\bm{0.827\pm 0.001}$ $0.801\pm 0.001$
jasmine $7.05$ $\mathit{0.831\pm 0.004}$ $\mathit{0.836\pm 0.003}$ $\bm{0.837\pm 0.003}$ $\mathit{0.835\pm 0.005}$
tecator $65.1$ $\mathit{0.969\pm 0.004}$ $0.959\pm 0.006$ $\bm{0.974\pm 0.003}$ $0.970\pm 0.004$

### v2

@@ Abstract
Some of the simplest, yet most frequently used predictors in statistics and machine learning use weighted linear combinations of features. Such linear predictors can model non-linear relationships between features by adding interaction terms corresponding to the products of all pairs of features. We consider the problem of accurately estimating coefficients for interaction terms in linear predictors. We hypothesize that the coefficients for different interaction terms have an approximate low-dimensional structure and represent each feature by a latent vector in a low-dimensional space. This low-dimensional representation can be viewed as a structured regularization approach that further mitigates overfitting in high-dimensional settings beyond standard regularizers such as the lasso and elastic net. We demonstrate that our approach, called LIT-LVM, achieves superior prediction accuracy compared to the elastic net, hierarchical lasso, and factorization machines on a wide variety of simulated and real data, particularly when the number of interaction terms is high compared to the number of samples. LIT-LVM also provides low-dimensional latent representations for features that are useful for visualizing and analyzing their relationships.
@@ 1 Introduction
Some of the simplest, yet most commonly used models in statistics and machine learning model an example’s target variable $y$ using a weighted linear combination of its features $x_{1},x_{2},\ldots,x_{p}$. These linear prediction models, such as linear regression and logistic regression, are favored for their speed, simplicity, and interpretability. In high-dimensional settings where the number of features $p$ is close to or even exceeds the number of examples $n$, regularization techniques such as the lasso Tibshirani (1996) and elastic net Zou & Hastie (2005), have become essential to prevent overfitting and enhance prediction accuracy on unseen test data.
@@
Unlike the original features, interaction terms have additional structure, as they denote relationships between two features. The coefficients for these interaction terms can be arranged in a $p\times p$ coefficient matrix $\mathbf{\bm{\Theta}}$, where entry $\theta_{jk}$ denotes the coefficient for the interaction term $x_{j}x_{k}$. Prior research on structured sparsity for interaction terms has typically assumed a hierarchical structure such that an interaction between features $j,k$ can only be present in the model if either feature $j$ or $k$ is in the model (Bien et al., 2013). We do not make such an assumption and further hypothesize that the coefficient matrix $\mathbf{\bm{\Theta}}$ has an approximate low-dimensional structure, so that we can represent each feature $j$ by a vector $\mathbf{\bm{z}}_{j}$ in a $d$-dimensional space, where $d<p$.
@@
Our main contribution is a structured regularization approach for estimating the coefficients for interaction terms in linear predictors using an approximate low-dimensional structure, as shown in Figure 1, in addition to traditional regularization techniques such as the elastic net. We call our proposed approach LIT-LVM, denoting linear and interaction terms with a latent variable model, and is applicable to many types of linear predictors, including linear regression, logistic regression, and the Cox proportional hazards (Cox PH) model (Cox, 1972). LIT-LVM fits in the space of models between elastic net with interactions, which provides sparse but unstructured interaction coefficient estimates; sparse factorization machines (Atarashi et al., 2021), which use exact low-rank factorizations of the interaction coefficients that are too rigid; and non-linear predictors, which are more expressive but not as interpretable as linear predictors with interactions.
@@ 2.1 Linear Predictors
[…]
@@ 2.2.1 Interaction Terms and Interpretability
Another line of research on interaction terms focuses on structured sparsity using the notion of heredity. Under the strong heredity structure, the interaction coefficient $\theta_{jk}\neq 0$ only if $\beta_{j}\neq 0$ and $\beta_{k}\neq 0$, implying that both the features $j,k$ must be selected in order to select their interaction. Several regularization methods that incorporate strong heredity or a relaxed version termed weak heredity have been proposed Zhao et al. (2009); Yuan et al. (2009); Choi et al. (2010); Radchenko & James (2010); Bien et al. (2013). The hierarchical lasso Bien et al. (2013) is probably the most similar to the methods we consider in this paper, as it also uses an elastic net penalty, where the $\ell_{1}$ portion enforces the strong or weak heredity structure. The heredity assumption simplifies the discovery of interaction terms, as one can first select a smaller number of features through their main effects $\mathbf{\bm{\beta}}$ and then consider only interactions between the selected terms, but does not assume any low-dimensional structure on the interaction matrix.
@@ 2.2.2 Latent Variable Models
2.2.3 Factorization Machines (FMs) and Related Models
The first FM model was proposed by Rendle (2010). It assumes that each feature $j$ has a $d$-dimensional latent vector $\mathbf{\bm{z}}_{j}$ such that $\mathbf{\bm{\Theta}}=\mathbf{\bm{Z}}\mathbf{\bm{Z}}^{T}$. This forces $\mathbf{\bm{\Theta}}$ to have an exact low-rank structure. Sparse FMs Xu et al. (2016); Pan et al. (2016); Atarashi et al. (2021) further constrain $\mathbf{\bm{\Theta}}$ to be sparse; we discuss these models further in Section 3.3.
Factorization of interaction terms has also been applied to models with time-varying coefficients through the factorized structured regression (FaStR) model proposed by Rügamer et al. (2022). FMs have been used for modeling in recommender systems, and the extension to time-varying coefficients in FaStR enables modeling in time-aware recommender systems where user behaviors change over time. FMs have also been extended to higher-order interaction terms in a scalable manner, including higher-order FMs Blondel et al. (2016) and additive higher-order FMs Rügamer (2024), which combine the benefits of GAMs with the scalability of factorized terms.
FMs and other models with factorized coefficients differ from our proposed LIT-LVM approach in two main aspects. First, they are heavily concerned with scalability when using interaction terms and are often applied to data with both large $n$ and $p$ with a sparse feature matrix $\mathbf{\bm{X}}$ Rendle (2012; 2013). On the other hand, our focus is on improving accuracy of estimation and prediction, not scalability, so our approach is not more scalable than directly modeling interaction terms. Second, FMs and other models with factorized coefficients assume that $\mathbf{\bm{\Theta}}$ is exactly low rank, whereas we assume that $\mathbf{\bm{\Theta}}$ has an approximate low-dimensional structure (possibly low rank, but not necessarily). This is a subtle, yet important distinction, and we provide more details in Section 3.3.
@@ 3.1 LVM for Interaction Terms
where the intercept $\alpha_{0}$ denotes the baseline interaction level. The latent distance model has a negative effect for distance so that features with more positive values of interactions are placed closer together in the latent space and more negative values of interactions are pushed farther apart. This type of negative effect on distance has also been used for matrices with both positive and negative entries by Huang & Xu (2022) and Nakis et al. (2023) and is often more interpretable than models with a positive effect on distance (Huang et al., 2022).
@@ 3.2 Estimation Procedure
We use Adam Kingma & Ba (2015) with various learning rates $\in$ $[0.005,0.01,0.05,0.1$] as our optimizer for the total loss function. To handle the $\ell_{1}$ penalty, we use the proximal gradient method Parikh & Boyd (2014), where the proximal operator corresponds to the soft thresholding operator. The other terms are all differentiable, with derivatives computed by automatic differentiation in PyTorch. The initializations of $\tilde{\mathbf{\bm{\beta}}}$ and $\mathbf{\bm{Z}}$ are drawn from normal distributions: $\tilde{\beta}_{j}\sim\mathcal{N}(0,1)$ and $z_{jk}\sim\mathcal{N}(0,1)$. The time complexity of our estimation procedure per epoch is $O(np^{2})$, as we discuss in Section A.2.
@@ 4 Simulation Experiments
We begin with some simulation experiments to investigate the effects of LIT-LVM as we change the number of features $p$ with fixed number of examples $n$. As $p$ increases, the model becomes more difficult to estimate. Code to reproduce all experiments is available at https://github.com/MLNS-Lab/LIT-LVM.
@@ Results
Figure 2: Results of simulation experiments on linear and logistic regression where the true latent dimension for the interaction weights $\mathbf{\bm{\Theta}}$ is $d_{\text{true}}=2$. (2(a)) Lower RMSE and (2(b)) higher AUC denote higher prediction accuracy. Error bars denote one standard error. Our proposed LIT-LVM approach, which uses structured regularization, significantly outperforms the elastic net with interaction terms, a traditional regularization approach, as $p$ increases for fixed $n$.
The results for fixed n=1,000n=1,\!000, true latent dimension $d_{\text{true}}=2$, and four different values of $p$ are shown in Figure 2(a). We compare our proposed LIT-LVM approach using a low rank model to elastic net with interaction terms. For $p=40$, $\mathbf{\bm{\Theta}}$ has $\binom{40}{2}=780$ entries, which is less than $n$, and both approaches are quite accurate. At $p=50$, $\mathbf{\bm{\Theta}}$ has $\binom{50}{2}=1,225$ entries, which now exceeds $n$, and we observe a phase transition where the RMSE of elastic net increases significantly compared to our LIT-LVM approach. As $p$ continues to increase, LIT-LVM continues to outperform the elastic net with interactions.
@@ 4.2 Logistic Regression with Latent Distance
In this experiment, we generate the feature matrix and interaction terms in the same manner as in the linear regression experiment in Section 4.1. For logistic regression, each element of the response vector $\mathbf{\bm{y}}$ is sampled from a Bernoulli distribution with success probability given by $\operatorname{logistic}(\mathbf{\bm{\beta}}^{T}\mathbf{\bm{x}}+\mathbf{\bm{\theta}}_{\text{flat}}^{T}\mathbf{\bm{x}}_{\text{int}})+\eta_{i}$, where $\operatorname{logistic}(x)=\frac{1}{1+e^{-x}}$. We inject sparsity into the augmented coefficient vector $\tilde{\mathbf{\bm{\beta}}}$ by first generating a dense version $\tilde{\mathbf{\bm{\beta}}}^{\text{dense}}$. We then set each element $\tilde{\beta}^{\text{dense}}_{j}$ to 0 with probability given by the Gaussian function $\exp(-(\tilde{\beta}^{\text{dense}}_{j})^{2}/\sigma_{s}^{2})$, where $\sigma_{s}^{2}$ is a parameter that adjusts the level of sparsity. This ensures that larger elements in $\tilde{\mathbf{\bm{\beta}}}^{\text{dense}}$ have a lower probability of being zeroed out.
@@ Results
The results for fixed n=1,000n=1,\!000, true latent dimension $d_{\text{true}}=2$, and five different values of $p$ are shown in Figure 2(b). Similar to the results from the linear regression experiment, we observe a phase transition where the accuracy of elastic net decreases significantly compared to our approach. The phase transition occurs at roughly $p=30$, lower than in the linear regression experiment. As $p$ continues to increase, our LIT-LVM approach continues to outperform elastic net, with prediction accuracy decreasing much slower than in elastic net.
@@
Comparison with Sparse FMs
In Section B.3, we conduct simulation experiments comparing the performance of LIT-LVM with sparse FMs [^1] To avoid differences in performance caused by the optimizer, we implement sparse FMs using LIT-LVM with $\lambda_{l}=100,000$. We show in Section C.3 that this value is sufficient. (Recall that LIT-LVM approaches a sparse FM in the limit as $\lambda_{l}\rightarrow\infty$.).. We simulate data in the same manner except with a low rank structure rather than a latent space structure, to better match the assumption of FMs. Likewise, we use LIT-LVM with a low rank model. We find that FMs perform worse when $\mathbf{\bm{\Theta}}$ is noisy or exhibits higher sparsity (deviation from exact low rank). This indicates that factorizing the interaction coefficients, as FMs do, can degrade performance when the underlying assumption of exact low rank structure is not met. In contrast, the approximate low rank structure assumption in LIT-LVM offers greater flexibility in modeling the interaction coefficients, ensuring more robustness and higher accuracy.
@@
Figure 3: Mean AUC for different values of the LVM regularization strength $\lambda_{l}$ on simulated data experiments using logistic regression. (3(a)) The simulation with low noise ($\sigma^{2}_{\theta}=0.1$) peaks in AUC at an intermediate $\lambda_{l}$, confirming the benefit of the LVM regularization term. (3(b)) The simulation with high noise ($\sigma^{2}_{\theta}=4$) has destroyed the latent structure, and thus shows minimal benefit from the LVM regularization.
@@
To assess the robustness of LIT-LVM under severe violations of the low-rank assumption, we repeat the same simulation but increase latent variable noise variance to $\sigma_{\theta}^{2}=4$, effectively destroying the underlying low-rank structure. As shown in Figure 3(b), the best performing $\lambda_{l}$ is near 0 or $0.01$ which corresponds to applying no or little LVM penalty. This is not surprising because at such high noise, the interaction matrix no longer possesses a low-rank structure so larger values of $\lambda_{l}$ degrade the performance.
@@ 5.1 OpenML Datasets
We conduct experiments on a wide variety of publicly available datasets from OpenML Vanschoren et al. (2013) to verify how well our structured regularization for the interaction matrix applies to real data. We consider both regression and binary classification tasks using linear and logistic regression, respectively. We consider 12 regression datasets and 10 classification datasets, described in Section C.1. For each task and dataset, we compare the accuracy of five different predictors, all with an elastic net penalty, listed below in increasing order of flexibility:
  1. 1.
    Elastic net with no interactions: the base linear predictor (linear or logistic regression) applied directly to the features.
  2. 2.
    Sparse factorization machines (FMs): factorized form for the interaction coefficient matrix that assumes exact low rank structure.
  3. 3.
    LIT-LVM: our proposed approach that assumes approximate low rank structure for the interaction coefficient matrix.
  4. 4.
    Hierarchical lasso: includes only interaction terms $x_{j}x_{k}$ where at least one of the features (main effects) $x_{j}$ or $x_{k}$ is also included. This corresponds to an assumption of weak heredity.
  5. 5.
    Elastic net with interactions: includes interaction terms between all pairs of features.
We divide the data into a 50/50 train/test split for both classification and regression tasks. We tune hyperparameters using a grid search, selecting the set of hyperparameters that minimize the mean RMSE for regression tasks and maximize the mean AUC for classification tasks over five splits within the training set. Subsequently, we evaluate prediction accuracy on the held-out test set. We repeat this process using five different train/test splits to calculate the standard errors for the prediction accuracy. Details on hyperparameter tuning are provided in Section C.1.2.
@@ Results
Dataset Name Elastic Net No Interactions Elastic Net with Interactions LIT-LVM Sparse FM Hierarchical Lasso
boston $4.99\pm 0.12$ $\mathbf{4.39\pm 0.23}$ $\mathit{4.40\pm 0.26}$ $\mathit{5.01\pm 0.63}$ $\mathit{4.51\pm 0.17}$
bike_sharing_demand $148.40\pm 1.08$ $\mathit{144.48\pm 1.06}$ $\mathbf{144.47\pm 1.05}$ $156.65\pm 1.07$ $\mathit{144.56\pm 0.39}$
diamonds $0.28\pm 0.00$ $0.26\pm 0.00$ $\mathbf{0.25\pm 0.00}$ $0.28\pm 0.00$ $0.28\pm 0.00$
fri_c4_500_10 $\mathit{0.88\pm 0.01}$ $\mathit{0.89\pm 0.02}$ $\mathbf{0.87\pm 0.02}$ $1.00\pm 0.03$ $\mathit{0.88\pm 0.02}$
medical_charges $0.24\pm 0.00$ $\mathbf{0.19\pm 0.00}$ $\mathbf{0.19\pm 0.00}$ $0.20\pm 0.00$ $0.20\pm 0.00$
nyc_taxi_green $0.50\pm 0.00$ $0.50\pm 0.00$ $\mathbf{0.16\pm 0.00}$ $0.50\pm 0.00$ $0.27\pm 0.00$
pol $30.55\pm 0.08$ $\mathit{24.33\pm 0.71}$ $\mathbf{24.24\pm 0.63}$ $29.05\pm 0.33$ $28.41\pm 0.06$
rmftsa_ladata $1.86\pm 0.05$ $\mathit{1.73\pm 0.05}$ $\mathbf{1.72\pm 0.03}$ $1.85\pm 0.05$ $6.49\pm 0.08$
socmob $23.57\pm 0.61$ $\mathit{17.17\pm 1.15}$ $\mathbf{16.55\pm 1.56}$ $23.56\pm 1.96$ $\mathit{17.52\pm 1.05}$
space_ga $0.14\pm 0.00$ $0.13\pm 0.00$ $\mathbf{0.12\pm 0.00}$ $0.13\pm 0.00$ $0.14\pm 0.00$
superconduct $17.72\pm 0.07$ $15.17\pm 0.19$ $\mathbf{14.32\pm 0.07}$ $17.70\pm 0.08$ $18.42\pm 0.05$
wind $3.30\pm 0.01$ $3.28\pm 0.03$ $\mathbf{3.17\pm 0.02}$ $3.28\pm 0.01$ $3.47\pm 0.02$
@@
Table 2: Mean AUC $\pm$ standard error over 5 different train/test splits for OpenML classification datasets. Higher AUC denotes higher prediction accuracy. Highest accuracy for each data set is denoted in bold, and entries within 1 standard error are denoted in italic. Our proposed LIT-LVM approach has the highest accuracy on 9 of the 10 datasets and is within 1 standard error on the 1 remaining dataset.
@@
Dataset Name Elastic Net No Interactions Elastic Net with Interactions LIT-LVM Sparse FM Hierarchical Lasso
Bioresponse $0.791\pm 0.013$ $0.791\pm 0.004$ $\bm{0.808\pm 0.003}$ $\mathit{0.801\pm 0.004}$ $\bm{0.808\pm 0.004}$
clean1 $0.900\pm 0.003$ $0.922\pm 0.004$ $\bm{0.947\pm 0.002}$ $0.900\pm 0.005$ $0.934\pm 0.003$
clean2 $0.944\pm 0.001$ $0.943\pm 0.003$ $\bm{0.985\pm 0.001}$ $0.955\pm 0.001$ $0.978\pm 0.001$
eye_movements $0.586\pm 0.004$ $\bm{0.608\pm 0.002}$ $\mathit{0.607\pm 0.002}$ $0.586\pm 0.004$ $\mathit{0.604\pm 0.005}$
fri_c4_500_100 $\bm{0.666\pm 0.011}$ $0.576\pm 0.020$ $\bm{0.666\pm 0.010}$ $\mathit{0.651\pm 0.012}$ $\bm{0.666\pm 0.010}$
fri_c4_1000_100 $\mathit{0.672\pm 0.009}$ $0.564\pm 0.009$ $\bm{0.689\pm 0.008}$ $\mathit{0.674\pm 0.014}$ $\mathit{0.678\pm 0.009}$
hill-valley $\mathit{0.711\pm 0.020}$ $0.547\pm 0.027$ $\bm{0.734\pm 0.034}$ $\mathit{0.699\pm 0.023}$ $\mathit{0.715\pm 0.020}$
jannis $0.814\pm 0.000$ $\mathit{0.826\pm 0.001}$ $\bm{0.827\pm 0.001}$ $0.801\pm 0.001$ $0.821\pm 0.001$
jasmine $\mathit{0.831\pm 0.004}$ $\mathit{0.836\pm 0.003}$ $\bm{0.837\pm 0.003}$ $\mathit{0.835\pm 0.005}$ $\mathit{0.832\pm 0.003}$
tecator $\mathit{0.969\pm 0.004}$ $0.959\pm 0.006$ $\bm{0.974\pm 0.003}$ $\mathit{0.970\pm 0.004}$ $\mathit{0.970\pm 0.003}$